Solve the system by elimination:
x-y=8
x-2y=10

Answer by stanbon(75887)

Solve the system by elimination:
x-y=8
x-2y=10
-----
Subtract the 2nd equation from the 1st and solve for "y":
y = -2
---
Substitute to solve for "x":
-2-y = 8
y = -2-8
y = -10
=============
Cheers,
Stan H.